OpenStudy (anonymous):

A bookstore sold 18,000 paperbacks one month. This was 10% less than the number of paperbacks the store sold the previous month. The following equation represents this situation, where x represents the number of paperbacks sold the previous month. x − 0.1x = 18,000 How many paperbacks did the store sell in both months combined? 1.) 20,000 2.) 35,990 3.) 38,000 4.) 180,000

jimthompson5910 (jim_thompson5910):

What do you get when you solve for x?

OpenStudy (alexwee123):

solve for x x − 0.1x = 18,000 0.9x=18,000 x=18000/0.9 x=?

OpenStudy (anonymous):

It's 38, 000! I figured it out! :D

jimthompson5910 (jim_thompson5910):

you got it, nice job